OpenStudy (lefinnois):

Am I right? What is the equation of the line that is parallel to y = –4x + 3 and has a y-intercept of -1/3?

OpenStudy (lefinnois):

Is it y=4x-1/3?

OpenStudy (adrimit):

almost, the slopes should be the same

OpenStudy (lefinnois):

Hmm.. This is multiple choice and they don't have -4x as an option.

OpenStudy (lefinnois):

They have 4x-y=1/3

OpenStudy (er.mohd.amir):

slope of parallel lines are same

OpenStudy (lefinnois):

Yeah, there isn't an option that's -4, unfortunately.

OpenStudy (adrimit):

ok if the lines are of the form ax+by=c, the slope is \(\frac{-a}{b}\)

OpenStudy (er.mohd.amir):

transfer -4x on other side

OpenStudy (lefinnois):

-4x=y-1/3?

OpenStudy (adrimit):

no, y+4x=-1/3

OpenStudy (er.mohd.amir):

y+4x=-1/3

OpenStudy (lefinnois):

Ah, thank you very much. :-)